# Trading Carbon for Physics: On the Resource Efficiency of Machine Learning for Spatio-Temporal Forecasting

### v1

@@ Appendix D Incompressible Shear Flow Dataset
The tracer field is initialized to match the shear flow, while the pressure field is initialized to zero everywhere in the domain. Thus, the initial conditions are fully characterized by three parameters: $n_{\text{shear}}\in\{2,4\}$, $n_{\text{blobs}}\in\{2,3,4,5\}$, and $w\in\{0.25,0.5,1.0,2.0,4.0\}$. An example of the data is illustrated in Fig. 9.
@@
As part of preprocessing, we standardize each field using $\hat{x}=(x-\mu)/\sigma$ (see Table 2). This normalization balances the contribution of different fields by preventing those with high variance from dominating the loss.
Table 2: Field statistics. Mean and standard deviation of raw fields, averaged over all time steps.
Field Mean Std
s $2.19\times 10^{-3}$ $0.33$
p $1.18\times 10^{-9}$ $0.074$
$\mathbf{v_{x}}$ $2.19\times 10^{-3}$ $0.40$
$\mathbf{v_{y}}$ $1.40\times 10^{-7}$ $0.097$
Appendix E Flow Matching in Low-Resolution Data Space
Most prior work applies FM in latent space, with a few exceptions (Disch et al., 2025; Baldan et al., 2025) demonstrating its feasibility directly in data space. We adopt this latter strategy to avoid the additional energy and carbon costs of an encoder–decoder pipeline. To make FM computationally tractable, we downsample inputs by a factor of 16, perform training and inference at low resolution, and subsequently upsample to the original resolution. This design inevitably encode a trade-off between efficacy and efficiency: while operating in low-resolution data space suppresses fine-scale dynamics and may introduce artifacts, it delivers substantial efficiency gains, reducing training and inference time by orders of magnitude.
Appendix F Model Configurations
@@
Model configurations are adapted from prior work. U-nets, FNO, and TFNO follow Ohana et al. (2024), UFNO follows Rahman et al. (2023), and FM is adapted from Lim et al. (2025), with some adjustments to match our setting (including operating directly in data space). In line with The Well (Ohana et al., 2024), U-nets and FNO variants are scaled to $\sim$17–20M parameters, while FM is smaller ($\sim$8.5M). Table 3 provides an overview of parameter counts, with detailed configurations summarized in Table 4 (U-nets and FNOs) and Table 5 (FM).
Table 3: Model parameters.
Model #Params
UNet 17.517.5M
UNet-CN 18.618.6M
UNet-CP 17.517.5M
UNet-CN-CP 18.618.6M
FM 8.68.6M
FNO 19.019.0M
TFNO 19.319.3M
UFNO 18.818.8M
Table 4: U-nets and FNOs model configurations.
@@
Model Filter size/modes Init dim Blocks/stage Up/down Bottleneck
UNet, UNet-CP 33 $48$ 11 44 11
UNet-CN, UNet-CN-CP 77 $42$ 22 44 11
FNO, TFNO $16$ $128$ 44
UFNO $8,12,16$ $128$ 11 33 11
Table 5: FM model configuration. The FM model operates in low-resolution data space and is built from transformer stacks with learned positional, timestamp, and distance embeddings.
Model Resolution Init dim Depth Bottleneck Heads
FM $16\times 32$ $256$ 33 44 88
Appendix G Additional Results
This appendix presents supplementary plots that support the findings in the main text. Table 6 reports both the training and inference emissions alongside run times.
Training Inference (one-step)
Model CO2eq (kg) Time (h) CO2eq (g) Time (min)
UNet $2.72$ $12.7$ $3.68\pm 0.03$ $1.0$
UNet-CN $4.71$ $15.9$ $7.77\pm 0.05$ $1.9$
UNet-CP $3.81$ $14.2$ $4.06\pm 0.05$ $1.0$
UNet-CN-CP $5.89$ $20.9$ $4.87\pm 0.02$ $1.2$
FM $3.31$ $14.0$ $23.54\pm 0.13$ $4.6$
FNO $0.80$ $4.0$ $7.87\pm 0.94$ $2.0$
TFNO $1.09$ $5.3$ $7.21\pm 0.12$ $1.8$
UFNO $1.12$ $3.6$ $9.00\pm 0.88$ $2.4$
Table 6: Carbon footprint during training and inference. Reported CO2eq emissions and runtime for all models, measured with CarbonTracker. Inference results correspond to one-step forecasts over the full test set (4,704 samples) and includes the initial model loading cost. Each inference experiment was repeated three times; CO2eq is reported as mean and standard deviation, and runtimes as mean values.
Fig.10-a reports model performance using the variance-scaled root mean squared error (VRMSE), which normalises prediction error by the variance of the ground truth to account for scale differences across fields:
$\text{VRMSE}(u,\hat{u})=\left(\frac{\langle|u-\hat{u}|^{2}\rangle}{\langle|u-\langle u\rangle|^{2}\rangle+\epsilon}\right)^{1/2},$ (19)
where $\langle\cdot\rangle$ denotes spatial averaging, $u$ the ground truth, $\hat{u}$ the prediction, and $\epsilon=10^{-8}$ a small constant for stability. The VRMSE measures errors relative to the natural variability of the data, with $\text{VRMSE}(y,\langle y\rangle)\approx 1$ implying that values greater than one indicate prediction errors larger than typical fluctuations. The results follow the same overall trends as in Fig.8-a, where performance was measured with the Pearson correlation coefficient.
Figure 10-b reports the inference carbon footprint for $24$ autoregressive rollouts of $50$ steps each. Compared to one-step inference over the full test set ($4,704$ samples), the rollout costs tend to average out across models, since the fixed overhead of model loading constitutes a larger share of the total cost. Flow matching remains an outlier, standing out as the most expensive model during inference.
Figure 10: Predictive performance and rollout carbon footprint. a: VRMSE vs. training carbon footprint. Each point shows every second step of a 20-step rollout, averaged across the four fields (lighter shades indicate earlier steps). b: Carbon footprint of inference during rollouts. Reported CO2eq emissions for 24 autoregressive rollouts of 50 steps. Inference costs include the initial model loading cost. Models are ordered from no bias (left) to strong bias (right).
Figs. 11, 12, and 13 provide rollout prediction examples for the three physical fields not shown in the main text (pressure fields and the two velocity components). The observations made for the tracer field in the main text extend to these additional fields.

### v2

@@ Appendix 0.B Incompressible Shear Flow Dataset
The tracer field is initialized to match the shear flow, while the pressure field is initialized to zero everywhere in the domain. Thus, the initial conditions are fully characterized by three parameters: $n_{\text{shear}}\in\{2,4\}$, $n_{\text{blobs}}\in\{2,3,4,5\}$, and $w\in\{0.25,0.5,1.0,2.0,4.0\}$.
@@
As part of preprocessing, we standardize each field using $\hat{x}=(x-\mu)/\sigma$. This normalization balances the contribution of different fields by preventing those with high variance from dominating the loss.
Appendix 0.C Model Configurations
@@
Model configurations are adapted from prior work. U-nets, FNO, and TFNO follow Ohana et al. [26], while UFNO follows Rahman et al. [30]. All models are scaled to 17–20M parameters. Table 3 summarises configurations and parameter counts.
Table 3: Model configurations.
@@
Model Filter size/modes Init dim Blocks/stage Up/down Bottleneck Params ($\times 10^{6}$)
UNet, UNet-CP 33 $48$ 11 44 11 $17.5$
UNet-CN, UNet-CN-CP 77 $42$ 22 44 11 $18.6$
FNO, TFNO $16$ $128$ 44 $19.0$, $19.3$
UFNO $8,12,16$ $128$ 11 33 11 $18.8$
Appendix 0.D Rollout Results
Figs. 8-11 provide rollout prediction examples for each of the four fields. The baseline U-net (UNet and UNet-CP) deteriorate rapidly, accumulating distortions within just a few steps. The modernised variants (UNet-CN, UNet-CN-CP) show marked improvements: UNet-CN captures fine-scale detail but gradually disrupts large-scale dynamics, while UNet-CN-CP better preserves structure at both scales at the cost of high-frequency artifacts. FNO and TFNO maintain dominant large-scale dynamics but exhibit a systematic drift and fail to capture fine-scale variability. UFNO recovers both large- and small-scale structures more effectively, though these gradually smooth out over time.
@@ Appendix 0.D Rollout Results
Figure 11: Example rollout predictions for the tracer field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).